OpenStudy (gabylovesyou):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (anonymous):

Sam is ten. His coach is 3x older than him, so his coach is 30. Five years ago, Sam was 5 and his coach was 25. 25/5=5

OpenStudy (gabylovesyou):

so the answer is 10

OpenStudy (anonymous):

@Snapbacklive how did you know sam was 10?

OpenStudy (anonymous):

Sam's coach is three times older than Sam' tells you y = 3x. Also, 5 years ago, Sam's age was x-5 and the coach's age was y-5. Now the coach's age was 5 times Sam's age, so y-5 = 5(x-5), and expanding out the right hand side y-5 = 5x - 25 Substituting y = 3x gives 3x - 5 = 5x - 25 Adding 25 to both sides 3x + 20 = 5x Subtracting 3x from both sides 2x = 20 Dividing through by 2, x = 10 and Sam is ten years old.

OpenStudy (anonymous):

@Gabylovesyou coach: 3x-5 = 5y....(i) sam: x-5 = y....(ii) substitute in (i) 3x-5 = 5(x-5) 3x-5 = 5x-25 3x-5x = -25+5 -2x = -20 x = 10

OpenStudy (anonymous):

yup

OpenStudy (gabylovesyou):

thank u!

OpenStudy (anonymous):

ur welcome